If on increasing the speed of a vehicle by 2m/s, the kinetic energy gets doubled, then what would be its initial speed?
1
Expert's answer
2018-03-19T08:09:08-0400
we got this equation:
E_2=2E_1
(m〖(ʋ_1+2)〗^2)/2= m〖ʋ_1〗^2
〖〖(ʋ〗_1+2)〗^2=2〖ʋ_1〗^2
〖ʋ_1〗^2-4ʋ_1-4=0
D=32
ʋ_1= (4± √32)/2
ʋ_(1 1)=-0.82 ʋ_(1 2)=4.82
Answer: ʋ1 = 4.82 m/s